OpenStudy (anonymous):

Find an equation of the line tangent to the graph of 6/sqrtx at the point (25,6/5) please help?

OpenStudy (mertsj):

Find the first derivative.

OpenStudy (anonymous):

I did find the first derivative then I plugged 25 into the f'(x) which will give me the slope. But that was wrong. I used y-y1 = m(x-x1)

OpenStudy (mertsj):

What did you get for the first derivative?

OpenStudy (anonymous):

after finding derivative whether by the quotient rule or just by re-organizing the equation, I get f'(x) = 6 * (-1/2)x^(-3/2) then, at x = 25, I got 0.024 = which is slope. y - y 1 = m (x-x1) I plugged 6/5 for y 1 = 0.024 for m and 25 for x1

OpenStudy (anonymous):

I got the equation 0024x+06 but still wrong?

OpenStudy (mertsj):

Probably should leave the slope in its exact form.

OpenStudy (mertsj):

Also, how is the slope not negative?

OpenStudy (mertsj):

\[f'(x)=\frac{-3}{x ^{\frac{3}{3}}}\]

OpenStudy (anonymous):

Wait let me try it again.

OpenStudy (anonymous):

Yea I got it right now because I didn't have minus

OpenStudy (anonymous):

Thanks!

OpenStudy (mertsj):

yw